OpenStudy (anonymous):

Solve 3/4(8p-20)=7 p=9 p=-9 p=35 p=-35

OpenStudy (dumbcow):

can you rewrite this please, not getting those answers

OpenStudy (anonymous):

I try to solve but notn macth

jhonyy9 (jhonyy9):

24p-60 =28 --- 24p=88 /8 --- 3p=11

OpenStudy (anonymous):

p=11/3....